OpenStudy (anonymous):

find the variance of 101, 102, 100, 100, 109, 110, 109, 108, 109

OpenStudy (anonymous):

Σ(x) = 101 + 102 + 100 + 100 + 110 + 109 + 109 + 108 + 109 = 948 Σ(x²) = 101² + 102² + 100² + 100² + 110² + 109² + 109² + 108² + 109² = 10201 + 10404 + 10000 + 10000 + 12100 + 11881 + 11881 + 11664 + 11881 = 100012 n = 9 Therefore Variance = 100012/9 - (948/9)² = 52/3 Standard deviation = √(σ²) = √(52/3) = 4.16(3sf)

OpenStudy (amistre64):

oyvey! id subtract like 105 from all of them and work the smaller set

OpenStudy (amistre64):

{101, 102, 100, 100, 109, 110, 109, 108, 109}- 105 {-4, -3, -5, -5, 4, 5, 4, 3, 4} thats almost mean centered already .... and alot less scary to work with

OpenStudy (amistre64):

sum {-4, -3, -5, -5, 4, 5, 4, 3, 4} = 3 mean = 3/9 = 1/3 ; the original sets mean is therefore 105 + 1/3 but the variance of each set is the same

OpenStudy (amistre64):

{-4, -3, -5, -5, 4, 5, 4, 3, 4} - 1/3 [ {-12, -9, -15, -15, 12, 15, 12, 9, 12} - 1]/3 [{-13, -10, -16, -16, 11, 14, 11, 8, 11}/3]^2 {169+100+256+256+121+196+121+64+121}/9 / 8 {169+100+256+256+121+196+121+64+121}/72 1404/72 = 39/2

OpenStudy (amistre64):

52/3 for a population :)